# TopoBench: A Framework for Benchmarking Topological Deep Learning

### v2

@@ A.1 Topological Domains
A topological domain is a generalization of a graph that captures both pairwise and higher-order relationships between entities (Bick et al., 2023; Battiston et al., 2021). When working with topological domains, two key properties come into play: set-type relations and hierarchical structures represented by rank functions (Hajij et al., 2023; Papillon et al., 2023).
@@
Simplicial complexes extend graphs by incorporating hierarchical part-whole relationships through the multi-scale construction of cells. In this structure, nodes correspond to rank 00-cells, which can be combined to form edges (rank $1$-cells). Edges can then be grouped to form faces (rank 2 cells), and faces can be combined to create volumes (rank $3$-cells), continuing in this manner. Consequently, the faces of a simplicial complex are triangles, volumes are tetrahedrons, and higher-dimensional cells follow the same pattern. A key feature of simplicial complexes is their strict hierarchical structure, where each $k$-dimensional simplex is composed of $(k-1)$-dimensional simplices, reinforcing a strong sense of hierarchy across all levels.
@@ A.2 Liftings
Lifting describes the process of mapping two topological domains through a well-defined procedure (Hajij et al., 2023; Papillon et al., 2023). This work extends this concept by providing a unified mathematical framework that generalizes all lifting procedures from the 2nd Topological Deep Learning Challenge at ICML 2024 (Bernárdez et al., 2024).
@@ A.2.1 Lifting Examples
From graphs to hypergraphs: $\boldsymbol{k}$-hop liftings. Let $\mathcal{G}=(V,E)$ be a graph and $\mathcal{H}=(V,\mathcal{E})$ be a hypergraph. The $k$-neighborhood $N_{k}(v)$ of a node $v\in V$ in $\mathcal{G}$ consists of all nodes reachable within $k$ steps from $v$. To lift $\mathcal{G}$ to $\mathcal{H}$, a hyperedge $e_{v}$ is assigned to each node $v\in V$ in $\mathcal{H}$, where $e_{v}=N_{k}(v)$. Thus, the set of hyperedges in $\mathcal{H}$ is given by =$\mathcal{E}\;=\;\bigl{\{}\,N_{k}(v)\,\bigm{|}\,v\in V\bigr{\}}$.
@@
Lifting maps can be either fixed (Bodnar et al., 2021a; Hajij et al., 2023) or learnable (Battiloro et al., 2024a; Bernárdez et al., 2023; Telyatnikov and Scardapane, 2023; Ramamurthy et al., 2023; Kazi et al., 2022), and they may compute or learn both the features on higher-order cells and the structure of the domain itself.
@@ A.3.1 General definition
Topological neural networks (TNNs) are neural architectures that process data defined on topological domains. The higher-order message passing paradigm of Hajij et al. (2023) provides a unifying framework for TNNs, and all networks used in TopoBench can be viewed as special cases of this approach.
@@ A.4 Traditional Message Passing on Graphs
Formally, a graph is defined as a tuple of nodes and edges, $\mathcal{G}=(\mathcal{V},\mathcal{E})$. In the context of k-cochain spaces introduced in Section A.3.1, we can view node features as 00-cochains and edge features as $1$-cochains. We denote by $h_{k}^{t}\in\mathbb{R}^{d}$ the hidden state of a node $k$ at MP iteration $t$, which can be seen as an element of a 00-cochain space.
@@ Appendix B Implemented Liftings
Table 3: List of implemented liftings, each one linked with its description. The domains are: PC for point clouds, G for graphs, HG for hypergraphs, SC for simplicial complexes, CC for cellular complexes, and CCC for combinatorial complexes.
@@
Clique lifting G SC
Neighborhood/dowker lifting G SC
Vietoris-rips lifting G SC
Graph induced lifting G SC
Line lifting G SC
Eccentricity lifting G SC
Dnd lifting G SC
Random latent clique lifting G SC
Neighbourhood complex lifting G SC
Alpha complex lifting PC SC
Random flag complex lifting PC SC
Cycle lifting G CC
Discrete conf. complex G CC
K-hop lifting G HG
Expander hypergraph lifting G HG
KNN-Lifting G HG
Kernel lifting G HG
Mapper lifting G HG
Modularity maximization lifting G HG
Forman-ricci curvature lifting G HG
Voronoi lifting PC HG
Pointnet++ lifting PC HG
@@
Simplicial paths lifting G CCC
Coface lifting SC CCC
Universal strict lifting HG CCC
@@ C.3 Descriptive Summaries of Datasets
Table 4 provides descriptive statistics for each dataset used in the numerical experiments (see Section 5 for details) after lifting to three topological domains: simplicial complexes, cellular complexes, and hypergraphs. The columns labeled 00-cell, $1$-cell, $2$-cell, and $3$-cell show the count of each $n$-cell in the resulting dataset. Specifically, a clique complex lifting is applied to obtain a simplicial domain with a maximum dimension of three, cycle-based lifting is used to obtain a cellular domain with a maximum dimension of two, and $k$-hop lifting (with $k=1$) is used to lift each graph into a hypergraph.
@@
Dataset Domain 00-cell $1$-cell $2$-cell $3$-cell Num. Hyperedges Cora Cellular 2708 5278 2648 0 0 Simplicial 2708 5278 1630 220 0 Hypergraph 2708 0 0 0 2708 Citeseer Cellular 3327 4552 1663 0 0 Simplicial 3327 4552 1167 255 0 Hypergraph 3327 0 0 0 3327 PubMed Cellular 19717 44324 23605 0 0 Simplicial 19717 44324 12520 3275 0 Hypergraph 19717 0 0 0 19717 MUTAG Cellular 3371 3721 538 0 0 Simplicial 3371 3721 0 0 0 Hypergraph 3371 0 0 0 3371 NCI1 Cellular 122747 132753 14885 0 0 Simplicial 122747 132753 186 0 0 Hypergraph 122747 0 0 0 122747 NCI109 Cellular 122494 132604 15042 0 0 Simplicial 122494 132604 183 0 0 Hypergraph 122494 0 0 0 122494 PROTEINS Cellular 43471 81044 38773 0 0 Simplicial 43471 81044 30501 3502 0 Hypergraph 43471 0 0 0 43471 REDDIT-BINARY Cellular 859254 995508 141218 0 0 Simplicial 859254 995508 49670 1303 0 Hypergraph 859254 0 0 0 859254 IMDB-BINARY Cellular 19773 96531 77758 0 0 Simplicial 19773 96531 391991 1694513 0 Hypergraph 19773 0 0 0 19773 IMDB-MULTI Cellular 19502 98903 80901 0 0 Simplicial 19502 98903 458850 2343676 0 Hypergraph 19502 0 0 0 19502 ZINC Cellular 277864 298985 33121 0 0 Simplicial 277864 298985 769 0 0 Hypergraph 277864 0 0 0 277864 Amazon Ratings Cellular 24492 93050 68553 0 0 Simplicial 24492 93050 110765 64195 0 Hypergraph 24492 0 0 0 24492 Minesweeper Cellular 10000 39402 28955 0 0 Simplicial 10000 39402 39204 9801 0 Hypergraph 10000 0 0 0 10000 Roman Empire Cellular 22662 32927 10266 0 0 Simplicial 22662 32927 7168 0 0 Hypergraph 22662 0 0 0 22662 Tolokers Cellular OOM OOM OOM OOM OOM Simplicial OOM OOM OOM OOM OOM Hypergraph 11758 0 0 0 11758 US-county-demos Cellular 3224 9483 6266 0 0 Simplicial 3224 9483 6490 225 0 Hypergraph 3224 0 0 0 3224
C.4 Additional Results
@@
Note that for demonstration purposes, only one fixed lifting is applied to transform graphs into each of the considered topological domains, leaving a domain-specific optimal lifting strategy beyond the scope of this paper. [^6] Learnable liftings may further optimize the predictive capacity of higher-order networks. Specifically, a clique complex is used for simplicial lifting, cycle-based lifting is used for cellular domains, and $k$-hop lifting (with $k=1$) is used for hypergraphs. Feature projection is also applied, where the $(n-1)$-cell features are multiplied by the corresponding incidence matrices to generate $n$-cell features.
@@
Node-level tasks Dataset GCN GIN GAT AST EDGNN UniGNN2 CCXN CWN CCCN SCCN SCCNN SCN Cora 87.09 ± 0.2 87.21 ± 1.89 86.71 ± 0.95 88.92 ± 0.44 87.06 ± 1.09 86.97 ± 0.88 86.79 ± 1.81 86.32 ± 1.38 87.44 ± 1.28 80.86 ± 2.16 82.19 ± 1.07 82.27 ± 1.34 Citeseer 75.53 ± 1.27 73.73 ± 1.23 74.41 ± 1.75 73.85 ± 2.21 74.93 ± 1.39 74.72 ± 1.08 74.67 ± 2.24 75.2 ± 1.82 75.63 ± 1.58 69.6 ± 1.83 70.23 ± 2.69 71.24 ± 1.68 Pubmed 89.4 ± 0.3 89.29 ± 0.41 89.44 ± 0.24 89.62 ± 0.25 89.04 ± 0.51 89.34 ± 0.45 88.91 ± 0.47 88.64 ± 0.36 88.52 ± 0.44 88.37 ± 0.48 88.18 ± 0.32 88.72 ± 0.5 Amazon 49.56 ± 0.55 49.16 ± 1.02 50.17 ± 0.59 50.5 ± 0.27 48.18 ± 0.09 49.06 ± 0.08 48.93 ± 0.14 51.9 ± 0.15 50.26 ± 0.17 OOM OOM OOM Empire 78.16 ± 0.32 79.56 ± 0.2 84.02 ± 0.51 79.5 ± 0.13 81.01 ± 0.24 77.06 ± 0.2 81.44 ± 0.31 81.81 ± 0.62 82.14 ± 0.0 88.27 ± 0.14 89.15 ± 0.32 88.79 ± 0.46 Minesweeper 87.52 ± 0.42 87.82 ± 0.34 89.64 ± 0.43 81.14 ± 0.05 84.52 ± 0.05 78.02 ± 0.0 $8$ 0.36 88.62 ± 0.04 89.42 ± 0.0 89.07 ± 0.25 89.0 ± 0.0 90.32 ± 0.11 Tolokers 83.02 ± 0.71 80.72 ± 1.19 84.43 ± 1.0 83.26 ± 0.1 77.53 ± 0.01 77.35 ± 0.03 OOM OOM OOM OOM OOM OOM Election 0.31 ± 0.02 0.28 ± 0.02 0.29 ± 0.02 0.29 ± 0.01 0.34 ± 0.02 0.37 ± 0.02 0.35 ± 0.02 0.34 ± 0.02 0.31 ± 0.02 0.53 ± 0.03 0.51 ± 0.03 0.46 ± 0.04 Bachelor 0.29 ± 0.02 0.31 ± 0.03 0.28 ± 0.02 0.3 ± 0.03 0.29 ± 0.02 0.31 ± 0.02 0.32 ± 0.03 0.33 ± 0.03 0.31 ± 0.02 0.36 ± 0.02 0.34 ± 0.03 0.32 ± 0.02 Birth 0.72 ± 0.09 0.72 ± 0.09 0.71 ± 0.09 0.71 ± 0.08 0.7 ± 0.07 0.73 ± 0.1 0.74 ± 0.11 0.72 ± 0.09 0.71 ± 0.09 0.82 ± 0.09 0.79 ± 0.12 0.71 ± 0.08 Death 0.51 ± 0.04 0.52 ± 0.04 0.51 ± 0.04 0.49 ± 0.05 0.52 ± 0.05 0.51 ± 0.05 0.54 ± 0.06 0.54 ± 0.06 0.54 ± 0.06 0.58 ± 0.06 0.55 ± 0.05 0.52 ± 0.05 Income 0.22 ± 0.03 0.21 ± 0.02 0.2 ± 0.02 0.21 ± 0.02 0.23 ± 0.02 0.23 ± 0.02 0.25 ± 0.03 0.25 ± 0.03 0.23 ± 0.02 0.29 ± 0.03 0.28 ± 0.03 0.25 ± 0.02 Migration 0.8 ± 0.12 0.8 ± 0.1 0.77 ± 0.13 0.78 ± 0.12 0.8 ± 0.12 0.79 ± 0.12 0.85 ± 0.18 0.84 ± 0.13 0.84 ± 0.12 0.91 ± 0.18 0.9 ± 0.14 0.92 ± 0.2 Unempl 0.25 ± 0.03 0.22 ± 0.02 0.23 ± 0.03 0.22 ± 0.02 0.26 ± 0.03 0.28 ± 0.02 0.27 ± 0.03 0.25 ± 0.03 0.24 ± 0.03 0.43 ± 0.04 0.43 ± 0.04 0.38 ± 0.04 Graph-level tasks MUTAG 69.79 ± 6.8 79.57 ± 6.13 72.77 ± 2.77 71.06 ± 6.49 80.0 ± 4.9 80.43 ± 4.09 74.89 ± 5.51 80.43 ± 1.78 77.02 ± 9.32 70.64 ± 5.9 76.17 ± 6.63 73.62 ± 6.13 PROTEINS 75.7 ± 2.14 75.2 ± 3.3 76.34 ± 1.66 76.63 ± 1.74 73.91 ± 4.39 75.2 ± 2.96 75.63 ± 2.57 76.13 ± 2.7 73.33 ± 2.3 75.05 ± 2.76 74.19 ± 2.86 75.27 ± 2.14 NCI1 72.86 ± 0.69 74.26 ± 0.96 75.0 ± 0.99 75.18 ± 1.24 73.97 ± 0.82 73.02 ± 0.92 74.86 ± 0.82 73.93 ± 1.87 76.67 ± 1.48 76.17 ± 1.39 76.6 ± 1.75 74.49 ± 1.03 NCI109 72.2 ± 1.22 74.42 ± 0.7 73.8 ± 0.73 73.75 ± 1.09 74.93 ± 2.5 70.76 ± 1.11 75.66 ± 1.3 73.8 ± 2.06 75.35 ± 1.5 75.49 ± 1.39 77.12 ± 1.07 75.7 ± 1.04 IMDB-BIN 72.0 ± 2.48 70.96 ± 1.93 69.76 ± 2.65 70.32 ± 3.27 69.12 ± 2.92 71.04 ± 1.31 70.08 ± 1.21 70.4 ± 2.02 69.12 ± 2.82 70.88 ± 3.98 70.88 ± 2.25 70.8 ± 2.38 IMDB-MUL 49.97 ± 2.16 47.68 ± 4.21 50.13 ± 3.87 50.51 ± 2.92 49.17 ± 4.35 49.76 ± 3.55 47.63 ± 3.45 49.71 ± 2.83 47.79 ± 3.45 49.71 ± 3.7 48.75 ± 3.98 49.49 ± 5.08 REDDIT 76.24 ± 0.54 81.96 ± 1.36 75.68 ± 1.0 74.84 ± 2.68 83.24 ± 1.45 75.56 ± 3.19 82.84 ± 2.54 85.52 ± 1.38 85.12 ± 1.29 74.44 ± 1.74 77.24 ± 1.87 71.28 ± 2.06 ZINC 0.62 ± 0.01 0.57 ± 0.04 0.61 ± 0.01 0.59 ± 0.02 0.51 ± 0.01 0.6 ± 0.01 0.4 ± 0.04 0.34 ± 0.01 0.34 ± 0.02 0.46 ± 0.08 0.36 ± 0.02 0.53 ± 0.04
@@
Table 6: An ablation study compares the performance of CCXN, CWN, CCCN, SCCN, SCCNN, and SCN models on various datasets using two readout strategies, direct readout (DR) and signal down-propagation (SDP). SDP generally enhances CWN performance, whereas the effect of SDP on CCCN, SCCNN, and SCN varies based on their internal signal propagation mechanisms. Means and standard deviations of performance metris are shown. The best results are shown in bold for each model and readout type.

### v3

@@ A.1 Topological Domains
A topological domain is a generalization of a graph that captures both pairwise and higher-order relationships between entities (Bick et al., 2023; Battiston et al., 2021). When working with topological domains, two key properties come into play: set-type relations and hierarchical structures represented by rank functions (Hajij et al., 2023b; Papillon et al., 2023).
@@
Simplicial complexes extend graphs by incorporating hierarchical part-whole relationships through the multi-scale construction of cells. In this structure, nodes correspond to rank 0-cells, which can be combined to form edges (rank $1$-cells). Edges can then be grouped to form faces (rank 2 cells), and faces can be combined to create volumes (rank $3$-cells), continuing in this manner. Consequently, the faces of a simplicial complex are triangles, volumes are tetrahedrons, and higher-dimensional cells follow the same pattern. A key feature of simplicial complexes is their strict hierarchical structure, where each $k$-dimensional simplex is composed of (k1)(k-1)-dimensional simplices, reinforcing a strong sense of hierarchy across all levels.
@@ A.2 Liftings
Lifting describes the process of mapping two topological domains through a well-defined procedure (Hajij et al., 2023b; Papillon et al., 2023). This work extends this concept by providing a unified mathematical framework that generalizes all lifting procedures from the 2nd Topological Deep Learning Challenge at ICML 2024 (Bernárdez et al., 2024).
@@ A.2.1 Lifting Examples
From graphs to hypergraphs: $\boldsymbol{k}$-hop liftings. Let $\mathcal{G}=(V,E)$ be a graph and $\mathcal{H}=(V,\mathcal{E})$ be a hypergraph. The $k$-neighborhood $N_{k}(v)$ of a node $v\in V$ in $\mathcal{G}$ consists of all nodes reachable within $k$ steps from $v$. To lift $\mathcal{G}$ to $\mathcal{H}$, a hyperedge $e_{v}$ is assigned to each node $v\in V$ in $\mathcal{H}$, where $e_{v}=N_{k}(v)$. Thus, the set of hyperedges in $\mathcal{H}$ is given by $\mathcal{E}\;=\;\bigl{\{}\,N_{k}(v)\,\bigm{|}\,v\in V\bigr{\}}$.
@@
Lifting maps can be either fixed (Bodnar et al., 2021a; Hajij et al., 2023b) or learnable (Battiloro et al., 2024a; Bernárdez et al., 2023; Telyatnikov and Scardapane, 2023; Ramamurthy et al., 2023; Kazi et al., 2022), and they may compute or learn both the features on higher-order cells and the structure of the domain itself.
@@ A.3.1 General definition
Topological neural networks (TNNs) are neural architectures that process data defined on topological domains. The higher-order message passing paradigm of Hajij et al. (2023b) provides a unifying framework for TNNs, and all networks used in TopoBench can be viewed as special cases of this approach.
@@ A.4 Traditional Message Passing on Graphs
Formally, a graph is defined as a tuple of nodes and edges, $\mathcal{G}=(\mathcal{V},\mathcal{E})$. In the context of k-cochain spaces introduced in Section A.3.1, we can view node features as 0-cochains and edge features as $1$-cochains. We denote by $h_{k}^{t}\in\mathbb{R}^{d}$ the hidden state of a node $k$ at MP iteration $t$, which can be seen as an element of a 0-cochain space.
@@ Appendix B Implemented Liftings
Table 3: List of implemented liftings, each one linked with its description. The domains are: PC for point clouds, G for graphs, HG for hypergraphs, SC for simplicial complexes, $C$ for cellular complexes, and CCC for combinatorial complexes.
@@
Clique G SC
Neighborhood G SC
Vietoris-Rips G SC
Graph induced G SC
Line G SC
Eccentricity G SC
DnD G SC
Random latent clique G SC
Neighborhood complex G SC
Alpha complex PC SC
Random flag complex PC SC
Cycle G $C$
Discrete configuration G $C$
K-hop G HG
Expander hypergraph G HG
KNN G HG
Kernel G HG
Mapper G HG
Modularity maximization G HG
Forman-Ricci curvature G HG
Voronoi PC HG
PointNet++ PC HG
@@
Simplicial paths G CCC
Coface SC CCC
Universal strict HG CCC
@@ C.3 Descriptive Summaries of Datasets
Table 4 provides descriptive statistics for each dataset used in the numerical experiments (see Section 5 for details) after lifting to three topological domains: simplicial complexes, cellular complexes, and hypergraphs. The columns labeled 0-cell, 11-cell, 22-cell, and $3$-cell show the count of each $n$-cell in the resulting dataset. Specifically, a clique complex lifting is applied to obtain a simplicial domain with a maximum dimension of three, cycle-based lifting is used to obtain a cellular domain with a maximum dimension of two, and $k$-hop lifting (with $k=1$) is used to lift each graph into a hypergraph.
@@
Dataset Domain 0-cell 11-cell 22-cell 33-cell Num. Hyperedges Cora Cellular 2708 5278 2648 0 0 Simplicial 2708 5278 1630 220 0 Hypergraph 2708 0 0 0 2708 Citeseer Cellular 3327 4552 1663 0 0 Simplicial 3327 4552 1167 255 0 Hypergraph 3327 0 0 0 3327 PubMed Cellular 19717 44324 23605 0 0 Simplicial 19717 44324 12520 3275 0 Hypergraph 19717 0 0 0 19717 MUTAG Cellular 3371 3721 538 0 0 Simplicial 3371 3721 0 0 0 Hypergraph 3371 0 0 0 3371 NCI1 Cellular 122747 132753 14885 0 0 Simplicial 122747 132753 186 0 0 Hypergraph 122747 0 0 0 122747 NCI109 Cellular 122494 132604 15042 0 0 Simplicial 122494 132604 183 0 0 Hypergraph 122494 0 0 0 122494 PROTEINS Cellular 43471 81044 38773 0 0 Simplicial 43471 81044 30501 3502 0 Hypergraph 43471 0 0 0 43471 REDDIT-BINARY Cellular 859254 995508 141218 0 0 Simplicial 859254 995508 49670 1303 0 Hypergraph 859254 0 0 0 859254 IMDB-BINARY Cellular 19773 96531 77758 0 0 Simplicial 19773 96531 391991 1694513 0 Hypergraph 19773 0 0 0 19773 IMDB-MULTI Cellular 19502 98903 80901 0 0 Simplicial 19502 98903 458850 2343676 0 Hypergraph 19502 0 0 0 19502 ZINC Cellular 277864 298985 33121 0 0 Simplicial 277864 298985 769 0 0 Hypergraph 277864 0 0 0 277864 Amazon Ratings Cellular 24492 93050 68553 0 0 Simplicial 24492 93050 110765 64195 0 Hypergraph 24492 0 0 0 24492 Minesweeper Cellular 10000 39402 28955 0 0 Simplicial 10000 39402 39204 9801 0 Hypergraph 10000 0 0 0 10000 Roman Empire Cellular 22662 32927 10266 0 0 Simplicial 22662 32927 7168 0 0 Hypergraph 22662 0 0 0 22662 Tolokers Cellular OOM OOM OOM OOM OOM Simplicial OOM OOM OOM OOM OOM Hypergraph 11758 0 0 0 11758 US-county-demos Cellular 3224 9483 6266 0 0 Simplicial 3224 9483 6490 225 0 Hypergraph 3224 0 0 0 3224
Table 5 reports additional descriptive statistics for the graph datasets used in the experiments. Specifically, the table includes the dimensionality of the number of classes (set to 1 for regression tasks and to the actual class count for classification tasks), and the number of graphs in each dataset and the initial 0-cell (node) features. Note that, for the illustrative experiments in Section 5, a projected sum feature lifting is applied. Consequently, higher-order cells (e.g., $1$-cells, 22-cells, etc.) inherit the same initial feature dimensionality as the 0-cells.
Table 5: Additional descriptive statistics of the graph datasets used in the experiments.
Dataset 0-cell dim Num. classes Num. graphs Graph Cora 1433 7 1 Citeseer 3703 6 1 PubMed 19717 500 3 MUTAG 7 2 188 NCI1 37 2 4110 NCI109 38 2 4127 PROTEINS 3 2 1113 REDDIT-BINARY 10 2 2000 IMDB-BINARY 136 2 1000 IMDB-MULTI 89 3 1500 ZINC 21 1 12000 Amazon Ratings 300 5 1 Minesweeper 7 2 1 Roman Empire 300 18 1 Tolokers 10 2 1 US-county-demos 6 1 1
C.4 Additional Results and Analysis
@@
Note that for demonstration purposes, only one fixed lifting is applied to transform graphs into each of the considered topological domains, leaving a domain-specific optimal lifting strategy beyond the scope of this paper. [^5] Learnable liftings may further optimize the predictive capacity of higher-order networks. Specifically, a clique complex is used for simplicial lifting, cycle-based lifting is used for cellular domains, and $k$-hop lifting (with $k=1$) is used for hypergraphs. Feature projection is also applied, where the (n1)(n-1)-cell features are multiplied by the corresponding incidence matrices to generate $n$-cell features.
Finally, Tables 8 and 9 present the number of learnable parameters for each best-performing model configuration and their corresponding runtimes. Overall, these results indicate that TNNs tend to be less efficient in terms of memory usage and computational time compared to their graph-based counterparts. However, there are exceptions: EDGNN and UniGNN2 achieve parameter counts comparable to graph baselines, and among the TNNs, AST and EDGNN stand out as the most efficient on average.
@@
Node-level tasks Dataset GCN GIN GAT AST EDGNN UniGNN2 CCXN CWN CCCN SCCN SCCNN SCN Cora 87.09 ± 0.2 87.21 ± 1.89 86.71 ± 0.95 88.92 ± 0.44 87.06 ± 1.09 86.97 ± 0.88 86.79 ± 1.81 86.32 ± 1.38 87.44 ± 1.28 80.86 ± 2.16 82.19 ± 1.07 82.27 ± 1.34 Citeseer 75.53 ± 1.27 73.73 ± 1.23 74.41 ± 1.75 73.85 ± 2.21 74.93 ± 1.39 74.72 ± 1.08 74.67 ± 2.24 75.2 ± 1.82 75.63 ± 1.58 69.6 ± 1.83 70.23 ± 2.69 71.24 ± 1.68 Pubmed 89.4 ± 0.3 89.29 ± 0.41 89.44 ± 0.24 89.62 ± 0.25 89.04 ± 0.51 89.34 ± 0.45 88.91 ± 0.47 88.64 ± 0.36 88.52 ± 0.44 88.37 ± 0.48 88.18 ± 0.32 88.72 ± 0.5 Amazon 49.56 ± 0.55 49.16 ± 1.02 50.17 ± 0.59 50.5 ± 0.27 48.18 ± 0.09 49.06 ± 0.08 48.93 ± 0.14 51.9 ± 0.15 50.26 ± 0.17 OOM OOM OOM Empire 78.16 ± 0.32 79.56 ± 0.2 84.02 ± 0.51 79.5 ± 0.13 81.01 ± 0.24 77.06 ± 0.2 81.44 ± 0.31 81.81 ± 0.62 82.14 ± 0.0 88.27 ± 0.14 89.15 ± 0.32 88.79 ± 0.46 Minesweeper 87.52 ± 0.42 87.82 ± 0.34 89.64 ± 0.43 81.14 ± 0.05 84.52 ± 0.05 78.02 ± 0.0 88.88 ± 0.36 88.62 ± 0.04 89.42 ± 0.0 89.07 ± 0.25 89.0 ± 0.0 90.32 ± 0.11 Tolokers 83.02 ± 0.71 80.72 ± 1.19 84.43 ± 1.0 83.26 ± 0.1 77.53 ± 0.01 77.35 ± 0.03 OOM OOM OOM OOM OOM OOM Election 0.31 ± 0.02 0.28 ± 0.02 0.29 ± 0.02 0.29 ± 0.01 0.34 ± 0.02 0.37 ± 0.02 0.35 ± 0.02 0.34 ± 0.02 0.31 ± 0.02 0.53 ± 0.03 0.51 ± 0.03 0.46 ± 0.04 Bachelor 0.29 ± 0.02 0.31 ± 0.03 0.28 ± 0.02 0.3 ± 0.03 0.29 ± 0.02 0.31 ± 0.02 0.32 ± 0.03 0.33 ± 0.03 0.31 ± 0.02 0.36 ± 0.02 0.34 ± 0.03 0.32 ± 0.02 Birth 0.72 ± 0.09 0.72 ± 0.09 0.71 ± 0.09 0.71 ± 0.08 0.7 ± 0.07 0.73 ± 0.1 0.74 ± 0.11 0.72 ± 0.09 0.71 ± 0.09 0.82 ± 0.09 0.79 ± 0.12 0.71 ± 0.08 Death 0.51 ± 0.04 0.52 ± 0.04 0.51 ± 0.04 0.49 ± 0.05 0.52 ± 0.05 0.51 ± 0.05 0.54 ± 0.06 0.54 ± 0.06 0.54 ± 0.06 0.58 ± 0.06 0.55 ± 0.05 0.52 ± 0.05 Income 0.22 ± 0.03 0.21 ± 0.02 0.2 ± 0.02 0.21 ± 0.02 0.23 ± 0.02 0.23 ± 0.02 0.25 ± 0.03 0.25 ± 0.03 0.23 ± 0.02 0.29 ± 0.03 0.28 ± 0.03 0.25 ± 0.02 Migration 0.8 ± 0.12 0.8 ± 0.1 0.77 ± 0.13 0.78 ± 0.12 0.8 ± 0.12 0.79 ± 0.12 0.85 ± 0.18 0.84 ± 0.13 0.84 ± 0.12 0.91 ± 0.18 0.9 ± 0.14 0.92 ± 0.2 Unempl 0.25 ± 0.03 0.22 ± 0.02 0.23 ± 0.03 0.22 ± 0.02 0.26 ± 0.03 0.28 ± 0.02 0.27 ± 0.03 0.25 ± 0.03 0.24 ± 0.03 0.43 ± 0.04 0.43 ± 0.04 0.38 ± 0.04 Graph-level tasks MUTAG 69.79 ± 6.8 79.57 ± 6.13 72.77 ± 2.77 71.06 ± 6.49 80.0 ± 4.9 80.43 ± 4.09 74.89 ± 5.51 80.43 ± 1.78 77.02 ± 9.32 70.64 ± 5.9 76.17 ± 6.63 73.62 ± 6.13 PROTEINS 75.7 ± 2.14 75.2 ± 3.3 76.34 ± 1.66 76.63 ± 1.74 73.91 ± 4.39 75.2 ± 2.96 75.63 ± 2.57 76.13 ± 2.7 73.33 ± 2.3 75.05 ± 2.76 74.19 ± 2.86 75.27 ± 2.14 NCI1 72.86 ± 0.69 74.26 ± 0.96 75.0 ± 0.99 75.18 ± 1.24 73.97 ± 0.82 73.02 ± 0.92 74.86 ± 0.82 73.93 ± 1.87 76.67 ± 1.48 76.17 ± 1.39 76.6 ± 1.75 74.49 ± 1.03 NCI109 72.2 ± 1.22 74.42 ± 0.7 73.8 ± 0.73 73.75 ± 1.09 74.93 ± 2.5 70.76 ± 1.11 75.66 ± 1.3 73.8 ± 2.06 75.35 ± 1.5 75.49 ± 1.39 77.12 ± 1.07 75.7 ± 1.04 IMDB-BIN 72.0 ± 2.48 70.96 ± 1.93 69.76 ± 2.65 70.32 ± 3.27 69.12 ± 2.92 71.04 ± 1.31 70.08 ± 1.21 70.4 ± 2.02 69.12 ± 2.82 70.88 ± 3.98 70.88 ± 2.25 70.8 ± 2.38 IMDB-MUL 49.97 ± 2.16 47.68 ± 4.21 50.13 ± 3.87 50.51 ± 2.92 49.17 ± 4.35 49.76 ± 3.55 47.63 ± 3.45 49.71 ± 2.83 47.79 ± 3.45 49.71 ± 3.7 48.75 ± 3.98 49.49 ± 5.08 REDDIT 76.24 ± 0.54 81.96 ± 1.36 75.68 ± 1.0 74.84 ± 2.68 83.24 ± 1.45 75.56 ± 3.19 82.84 ± 2.54 85.52 ± 1.38 85.12 ± 1.29 74.44 ± 1.74 77.24 ± 1.87 71.28 ± 2.06 ZINC 0.62 ± 0.01 0.57 ± 0.04 0.61 ± 0.01 0.59 ± 0.02 0.51 ± 0.01 0.6 ± 0.01 0.4 ± 0.04 0.34 ± 0.01 0.34 ± 0.02 0.46 ± 0.08 0.36 ± 0.02 0.53 ± 0.04
@@
Table 7: Ablation study comparing the performance of CCXN, CWN, CCCN, SCCN, SCCNN, and SCN models on various datasets using two readout strategies, direct readout (DR) and signal down-propagation (SDP). SDP generally enhances CWN performance, whereas the effect of SDP on CCCN, SCCNN, and SCN varies based on their internal signal propagation mechanisms. Means and standard deviations of performance metris are shown. The best results are shown in bold for each model and readout type.